NAOJ GW Elog Logbook 3.2

In the past days we tried to characterize the locking loop of the filter.
The loop transfer function for the filter cavity (sketched in figure1) is compose by different blocks
- G1 [Hz/V] = piezo actuator
- G2 [Hz/Hz] = SHG
- G3 [Hz/W] = cavity
- G4 [W/V] = photodiode + demodulation
- H [V/V]= servo
- inject noise on perturb
- measure EPS1/EPS2
- inject noise on perturb
- measure piezo mon/EPS2
- inject noise on RAMP
- measure EPS2/piezo mon
The amplitude of the loop transfer functions plotted so far are actually the square of the real amplitude. The problem comes from the way I treated data saved by the spectrum analyzer. Each file is composed of 3 columns: frequency, real part (a) and imaginary part (b) of the TF. Of course amplitude and phase are recovered by doing:
Amplitude = sqrt (a^2 +b^2)
Phase = angle (a+i*b)
Due to an oversight, I had replaced the sqare root with the absolute value in the amplitude computation. This explain the unexpected behaviour (1/f^2 instead of 1/f) of the openloop TF around the UGF.
We will upload soon new TFs measurements (taken by Yuefan and Marc on monday night) properly plotted.

The activity of locking characterization of the past days pointed out some issues that are worth to be reported.
open loop (urad) | closed loop (urad) | |
PR yaw | 1-2 | 0.5-1 |
PR pitch | 4-8 | 2-3 |
BS yaw | 3-4 | 2-3 |
BS pitch | 5-6 | 4-5 |
INPUT yaw | 2-3 | 1-2 |
INPUT pitch | 4-5 | 3-4 |
END yaw | 2-3 | 1-2 |
END pitch | 4-5 | 3-4 |

After a long alignment work we were able to make the IR and the green beam flashing in the cavity at the same time. In this video are shown superposed flashes in trasmission both of the green and the IR (the green beam as been cut at the second 5).
In this configuration we were able to lock the cavity on the green TEM00 but we coudn't check the IR condition since there was too much green light transmitted by the dichroic mirror before the camera installed on the end bench.
After solving this issue and installing the AOM we will start to look for the frequency shift needed to have both the IR and the green resonant on the fundamental mode.

Then we powered up the AOM, there was the other orders appearing, the one close to the 0 order is the 1st order which we need. We put a PD at the 1st order, connected it to the oscilloscope and started to turn the AOM to have maximum power in the 1st order. After finding the best position of the AOM, we tried to change the RF power we were sending to increase the power more. But when we tried to do this, we found out the power fluctuation is about 40-50% percent, it was much more than the tolerance.
We put PD both in the infrared path and the green path to check the fluctuation of the green is coming from the infrared or not. Then we found out since the PD aperture is quite small, so it will increase the degree of uncertainty we saw. So we changed the PD to the powermeter which has larger aperture. This time the infrared is quite stable and the green is fluctuating but much less than before. We guessed that the fluctuation of the green is coming from the changing of the SHG cavity alignment when people moving around or touching the bench. We are going the install the MZ and stabilize the green.
Then other thing we checked is that the SHG loop gain we are using now is 50. But with this gain,if we changed the power sent to the PZT, the cavity cannot pull the error signal back to zero when the cavity is locked. If we increase the gain, until it reaches 10000, there will be no oscillation appear, the cavity will be locked more stable, but one problem is that it will be quite easy also lock on the other mode. We will try to find a the optimal gain considering also the filter cavity.

- In order to increase the pump power safely, we cover the laser path with aluminum black walls and black paper.
- We also cover the probe part in order to stop air flow from hepa filters which is a possible noise source. Kuroki-san helped me on this tasks.
- I put a small translation stage below the half ball in the imaging unit in order to adjust better the alignment and I tried many times the alignment until I found a good signal for the surface reference sample, comparable with the one of last year. Also the bulk reference sample gives almost the same signal as last year.
- The parts for the 1310nm probe laser were delivered, I glued the golden prism mirror on a half inch post, I'm waiting for the glue to cure under the neon lamps



The PDH filter cavity signal has been calibrated injecting a line at 28 kHz (above the ugf ~ 10 kHz of the loop) on the “ramp” input of the electronic servo. The ramp input is summed to the PZT correction signal.
The amplitude of the 28 kHz line in Hz is obtained using the formula:
S_Hz = V_RMS (V) * sqrt(2)*100*2e6 Hz/V = 1.25e-6*sqrt(2)*100*2e6 = 353 Hz
Where V_RMS is the line amplitude measured by the Agilent spectrum analyser. The factor sqrt(2) is obtained to pass from the V_RMS to the line amplitude (the factor has been also experimentally verified looking the same line with the spectrum analyser and the oscilloscope).
The factor 100 is the reduction of the PZT_moni output . 2e6 Hz/V is the calibration of the PZT after the SHG.
Measuring the line at 28 kHz in the error signal and compensating for the cavity frequency pole is it possible to find the calibration factor K in V/Hz. The formula used is :
S_V = K(V/Hz)* S_Hz /sqrt(1+ (f/f_0)^2)
where f_0 = 1.5 kHz and S_V = sqrt(2)*38.9e-3 V
--> K = 2.9e-3 V/Hz
which seems to be in agreement with the calibration obtained looking the PDH signal when the cavity is freely swinging. In that case we see a peak-to-peak of the PDH of ~ 4 V for 1.5 kHz of the cavity line which correspond to a K = 2.7e-3 V/Hz. Note that when the cavity is freely swining we have also rining effects which can perturb this measurement.
We have also checked that reducing the frequency of the line sent to the PZT (with the same amplitude) to 14 kHz, the amplitude of the line of the error signal is multiplied by 2, as expected given the cavity pole. A more quantitative analysis (fully taking in account the effect of the loop) is necessary to check the position of the cavity pole.
Another test was to increase the amplitude of the line by a factor 10, thus having a 29 kHz line with amplitude of 3 kHz (two times the cavity width of the cavity). The cavity stays locked and the calibration factor measured is the same with the one measured with the line with an amplitude of 300 Hz. Increasing further the amplitude of the 28 kHz line to ~ 7 kHz (4 times the cavity linewidth) makes the lock more fragile, and sometimes the cavity unlocks. Moreover, an oscillation with a frequency of ~ 1 Hz appears in the error signal (but it is not accompanied with a similar oscillation in the transmitted power).

We observe that the PDH filter cavity signal has an offset of ~ 170 mV. See picture.
The offset is present even when the 78 MHz signal sent to the EO modulator is swithed off (and the 78 MHz sent to the local oscillator is ON). When both signals are OFF, we see a slowly varying offset between 200 mV and -200 mV, which also have an higher frequency oscillation. To be investigated.

We have measured the spectrum of the PZT correction signal sent to the laser when the cavity is locked, using the output PZT_mon (1/100 of the PZT correction signal). The spectrum is in the attached plot.Since in this region the gain of the loop is very high, the signal is proportional to the cavity length/frequency noise.
The calibration is 1 MHz/V (given by the manufacturer).
at 100 Hz we have ~ 700 nV/sqrt(Hz) corresponding to 70 Hz/sqrt(Hz), at 1 kHZ we have 100 nV/sqrt(Hz) corresponding to 10 Hz/sqrt(Hz)
The shape of the spectrum is compatible with the free running laser noise ~ 7-10 kHz /f Hz/sqrt(Hz) up to a few kHz. According to aother measurement, after ~4 kHz the spectrum is limited by a flat noise, which is compatible with the noise of the 100 kOhm resistor at the output of the PZT_moni signal. For f<10 Hz probably the mirror control noise and the seismic noise are limiting the spectrum.
We also see several 50 Hz harmonics. It is not clear if this harmonics can be reduced rearranging the grounds and if they have an impact on the RMS of the error signal of the filter cavity locked. To be investigated.

Summary of yesterday night work (thu 29-->fri 30). The goal was to make a characterization campaign for the cavity lock, in order to make it more stable.
1) Beam stability
In the past we observed an evident jitter of the beam. From a comparison of the spectra we were convinced that this was caused by the residual motion of BS and PR. In the past days we where able to improve the stability by improving the local control filters (a dedicated entry will follow).
We observed that the beam direction (observed by misaligning the input mirror) was drifting and we decided to test a new strategy to keep the mirror position. We change the local control filters in order to avoid to gain at low frequency (we changed a pole at 0.1 Hz with a double zero at 0.1 Hz and we controlled the mirror position not by adding an offset of the loop but simply sending a DC signal to the coils.
We coudn't see a major improvement in the performances.
We also observed the intermittence presence of spikes in the error signals from BS and PR which makes difficult to keep the cavity alignment.
Eventually the old controls (with integrators at low frequency) were restored.
2) Laser servo gain transfer function
We have set the gain of the servo in order to have ~10 kHz bandwidth. See the transfer function in fig.1. (in 1/f^4 mode)
At a first look, the TF behaves as expected. The data have been stored in the floppy disk and they will be compared with the model. The phase margin at ~10 kHz is about 40 degrees.
The transfer function has been measured with the Agilent 35670A spectrum analyser, with a swept sine with 50 mV ptp.
3) Servo parameters
- modulation depth = 1 V pp at 78 MHz (reduced with respect to before). This should correspond to a modulation depth of m= 0.185 rad.
- LO = 8.5 Vpp at 78 MHz (increased with respect to before)
- Demodulation phase = 111 deg
--> With this data the error signal is 3-4 V ptp, for a transmited signal of ~ 3-4 V depending on the alignment of the cavity (note that we did not checked the green laser power yesterday night)
- attenuation of the input signal =9.1
- PZT gain = 0.7
- thermal control gain = 3
- Threshold on the transmitted signal ~ 2 V
4) Auto-relock
With this configuration the cavity automatically locks when the transmitted power crosses the resonance. When the cavity unlocks, it relocks automatically. Note that the servo is always in the 1/f^4 configuration. The video shows the cavity locked, then the input mirror is on purpose misaligned, then it is re-aligned and the cavity re-locks.
5) Stability
During yesterday night lock the cavity was very stable. The plots 2 and 3 show the transmitted power (in cyan) and the error signal (in yellow) for 500 s. No actions were performed to realign the cavity on the second plot. Max transmitted power was ~ 4 V.

On Tuesday 27th june we managed to lock the laser on the filter cavity length.
In the first attachment there is a plot of the transmitted power during the lock acquisition, in the second there is a picture of the transmitted beam when the cavity is locked. A short movie of the the lock acquisition can be seen here.

Friday, June 23, 2017
Following the following procedure, I aligned the Imaging Unit for the HeNe probe beam.
1) move IU micrometer closer to the end of translation which would give you enough translation range to move the whole IU farther away in case you test thick objects;
That will complete rough alignment of the IU. The fine tuning is done by maximizing AC signal coming from the surface calibration piece. For that, try different micrometer positions around one you started with. For every position you have to center the probe (maximize the DC), maximize AC if needed. The maximum R should be close to the original R for the surface calibration. Then make scan with the bulk calibration piece.
According to the theory the signal is maximum when the detector at the Rayleigh length of the perturbation, experimentally we can check this changing the position of the blade and aligning again the imaging unit and measuring the signal. So, in order to maximize the signal, I repeated the procedure changing the position of the blade from 18mm to 12mm and 6mm but I got a lower signal, so I aligned it back to 18mm.
The absorption signal of the reference is similar to original value (the one we had since we bought the system) even if I changed by the 20% the waists of pump and probe,
Parameters: LD current = 0.8A, power without sample = 33mW

We changed the layout of the end bench, added one more mirror after the beam splitter to divide the beam into two, one is received by a screen and a camera was set to look at this screen, the other goes directly into the CCD. The mirror we used has the maximum reflection when it is putting in 45 degree, so we turned it a bit in order to have large enough power also for the transmission. The one received by the screen we took it as a reference for the alignment of the cavity and the CCD is to find a good mode matching. In this configuration we will need three images sending back from the end room at the same time(two on the end bench and one for the second target), the electrical board we are using now for the video only has two channels. So we took the same board from the west end, but it seems one channel of that board is broken, but it is enough for our current need. The fiber used to send the video signal now is '1-13,1-14,1-15,1-16', each board need two fibers. Also each monitor only can receive two channels, so we also took the monitor at the second target of west end to the central room.
2.Analog signal board
On the reflection path of the beam splitter on the end bench we put a PD before and use the receiver box to send this signal to central room for locking the cavity. But since the the aperture of the PD is limited, the spectrum we saw will also be effected by the alignment of the beam. So yesterday we changed the PD to PSD in order to have the information of the beam position. But in this way, we need to send back three analog signals together. We found two board for this and one of them have four channels. We tested the board with a sine wave sent from the end room board, and received it in the central room. The fiber they used for this board is too short to connect into the rack('1-11,1-12'), but the fiber system of TAMA is too complicated, so we just simply changed with longer spare fibers, now we are using '3-13,3-14' for this board.
From the board it seems we can change the gain and offset of each channel, but when we sent a 4V peak to peak signal, it pretty hard to see if it is changed or not, so maybe we can only changed in very small scale. Also it seems four channels all have different setting offset and gains, but we need further check about this.

So we decided to increase laser current to have higher power in infrared and also in green, the current we used before is 1.040A which gave a 8mW-green power at the end of the bench, now we increased the current to 1.2A which gives out a three times green power around 27mW at the end of the bench.
With this power we easily found the beam at the second target and also received it at the end bench.By looking through the window of the end chamber, we better adjust the beam on the end mirror and let it pass more or less the center of it, luckily we got the flash of the cavity with the first try.
The other thing we did before closing the chamber is that we sent the beam out of the chamber to the corridor again after the BS, tried to superpose the green and infrared both in the near field and the far field with the last two infrared steering mirror on the bench. Although considering the air fluctuation in the corridor, two beams moved a lot, but we did our best to make them overlap at 300m.

2-inch mirror: f=-30cm z=4.51m
PR suspension: f=3m z=7.235m
Input mirror: f=-218.35m z=12.135m
End mirror: f=-218.35m z=312.135m
With these values, the simulation result shows that the beam will keep diverging in the arm which is not what we want(pic1). So I did some calculation, we need to move the 2-inch 3.9cm backward to get the beam waist around 150m inside the arm. (pic2)
The other reason why we should move 2-inch mirror is after we installed the real input mirror, we got green reflect back from the from surface of that mirror. But it seems the reflected beam is much larger the original one.
Before moving the mirror, we set two reference point, one is the position of the beam in front of the input mirror, the other is the beam reflected back to the bench. Also after the input mirror we put two aluminum mirror to reflect the beam into the corridor to check the beam propagation. At first we only tried to move the picomotor of the 2-inch, but even we when we finished all the range, nothing changed. So we moved by hand little by little. Every time we moved, we recovered the beam with pitch and yaw of the 2-inch on two reference point and check the beam size at the entrance of the corridor and bench. After moving about 3cm, the reflected beam size seems fine, so we followed the beam inside the corridor, found it has beam waist around 150m, and the size is from 1cm to 2cm which is acceptable.
Then we took the reference of the BS chamber like what we did to the PR, with two aluminum mirror, the beam transmitted by the 2-inch and the BS suspension have been reflect out of the chamber.

The filter's fan part was fine, so it was enough to buy only the filter part. New HEPA filters were delivered to Tama last friday. I washed the prefilters with water, I cleaned the fans, replaced the filters and placed them back to the top of the absorption bench clean booth.

After 30'-45' the pressure was at 100 mTor so I started the turbo pump.
On Friday evening (June 17th) the pressure in the EM2 chamber reached 3.1e-7 Torr.
On the other side of the gate valve the pressure was 1.1 e-7 Torr.
So I open the large gate valve and immediately after the valve between the turbo pump and the tube
(the valve between the turbo pump and the EM2 chamber was already open).
The pressure went up a little bit in the 1e-6 Torr range and then came down again.
This morning (Monday June the 19th) the pressures are:
EM2 chamber = 1.6e-7 Torr
EM2 tube = 2.3e-7 Torr
Middle of the tube: 4.7e-8 Torr
Middle of the tube: 2.7e-8 Torr
NM2 tube = 8.6e-8 Torr (this sensor was not working but is working again now)
NM2 chamber = in air
Overall it looks reasonable to me even if there is probably some out-gassing coming
from the EM2 chamber.

Setting the limits at the translation stage generated a bug in the part of the VI where it sends the positioning commands: if the position exceeds the axis range, the VI goes in loop. I fixed it making a subVI "Read limits axis.vi" that reads the limits from the translation stage controller and set the minimum and maximum positions in the property node of the position controls of the VI.

The origin was set at the front surface of the beam splitter. The beam waist size is w0=51.2um and waist position is z=-9.5cm.
Then I took this as the initial value to calculate the lens we are going to use for the telescope. So the condition is, the MC input mirror is 60cm away from the origin, and the beam waist should be around the position of the input mirror and has a size of 277um.
I let the program chose the lenses with focal length we have now and got this result.
L1: f1=100mm z1=5cm
L2: f2=200mm z2=37.5cm
L3: f3=200mm z3=55cm
So with this design the final beam waist is w1=279um, z1=60cm, which is the result we want.
The other thing is before we were using a 200mm lens to focalize the green beam we distracted from the Faraday to the PD. But yesterday we found out, if we want to give power supply and the signal output, we need more space for cable. But the position of the PD before was too close to the post of the steering mirror. So today I changed the lens into a 150mm one, and moved the PD 5cm forward.